# Search for joint multimessenger signals from potential Galactic PeVatrons with HAWC and IceCube

### v1

@@ 4.1 HAWC best-fit result
First, we present the best-fit results for the gamma-ray emission from the HAWC Pass 5 data alone. Table 2 summarizes the best-fit location, extension, and spectral parameters of each source in our list. In addition, we report the statistical and systematic uncertainty of the spectral fit results. Systematic uncertainties of the HAWC detector include PMT efficiency, charge uncertainty, etc., and are described in Abeysekara et al. (2019b). To evaluate their effect, we perform the model selection using samples simulated with different settings and compare the results with the baseline simulation. The errors of each type of systematic uncertainty are added in quadrature to obtain the total systematic uncertainty, which is given in the table 2.
| Name | RA [∘] | Dec [∘] | Ext [∘] | [TeV-1cm-2s-1] | Nearest TeVCat source | ||
| 3HWC J0534+220 | 83.64 | 22.01 | PS | 2.82 | 0.12 | Crab | |
| 3HWC J0634+067 | 98.59 | 6.66 | 0.49 | 2.57 | 0 | HAWC J0635+070 | |
| 3HWC J1757240 | 269.32 | 23.86 | PS | 2.58 | 0 | HESS J1800240B | |
| 3HWC J1809190 | 272.42 | 19.34 | 0.23 | 2.09 | 0.23 | HESS J1809193 | |
| 3HWC J1813125 | 273.4 | 12.7 | 0.27 | 2.6 | 0 | HESS J1813-126 | |
| 3HWC J1813174 | 273.4 | 17.74 | 0.26 | 1.91 | 0.35 | 2HWC J1814173 | |
| 3HWC J1825134 | 276.49 | 13.63 | 0.46 | 2.45 | 0.13 | 2HWC J1825134 | |
| 3HWC J1831095 | 277.85 | 9.85 | 1.14 | 2.59 | 0.14 | HESS J1831098 | |
| 3HWC J1837066 | 279.38 | 6.83 | 0.65 | 2.69 | 0.1 | 2HWC J1837065 | |
| 3HWC J1843034 | 280.95 | 3.39 | 0.6 | 2.51 | 0.13 | 2HWC J1844032 | |
| 3HWC J1847017 | 282.07 | 1.78 | 0.75 | 2.66 | 0.09 | HESS J1848018 | |
| 3HWC J1849+001 | 282.31 | 0.02 | 0.68 | 2.42 | 0.13 | IGR J184900000 | |
| 3HWC J1857+027 | 284.35 | 2.82 | 0.6 | 2.73 | 0.13 | HESS J1857+026 | |
| 3HWC J1908+063 | 287.02 | 6.35 | 0.57 | 2.44 | 0.09 | MGRO J1908+06 | |
| 3HWC J1912+103 | 288.1 | 10.32 | 0.52 | 2.83 | 0.14 | HESS J1912+101 | |
| 3HWC J1914+118 | 288.71 | 11.79 | 0.71 | 2.69 | 0 | 2HWC J1914+117* | |
| 3HWC J1922+140 | 290.74 | 14.06 | 0.13 | 2.77 | 0 | W 51 | |
| 3HWC J1928+178 | 292.13 | 17.81 | 0.83 | 2.52 | 0.11 | 2HWC J1928+177 | |
| 3HWC J1930+188 | 292.56 | 18.81 | 0.76 | 2.51 | 0.14 | SNR G054.1+00.3 | |
| 3HWC J2019+367 | 304.9 | 36.77 | 0.32 | 2.04 | 0.31 | VER J2019+368 | |
| 3HWC J2031+415 | 308.01 | 41.49 | 0.51 | 2.52 | 0.19 | TeV J2032+4130 | |
| 3HWC J2227+610 | 336.82 | 60.94 | PS | 2.42 | 0 | Boomerang |
@@ 4.2 Joint-fit result
We performed the joint fit using the best-fit model of HAWC and added a neutrino source as described in Section 3.4. We found no significant neutrino emission. The smallest pre-trial p-value from the individual source search is 0.07 for 3HWC J2019+367, which corresponds to a post-trial p-value of 0.21 when accounting for the trials in the individual source search. The pre-trial p-value for each source can be found in appendix B. We performed the binomial test and found no significant neutrino emission from any subset of the 22 sources. The pre-trial p-value of the binomial test is 0.09 at k=6 (3HWC J2019+367, 3HWC J2031+415, 3HWC J1912+103, 3HWC J1908+063, 3HWC J1857+027, 3HWC J1928+178) and corresponds to a post-trial p-value of 0.34 when only accounting for trials in the binomial test. Accounting for both the trials from the individual source search and the binomial test, the final post-trial p-value is 0.37.
@@
Based on the null result of the joint search, we also calculated the neutrino flux limits at 90% CL for the 22 sources (see table 3). We account for IceCube systematics by calculating the sensitivity using different Monte Carlo systematic sets and adding the difference to the baseline in quadrature. Appendix C details the effect of IceCube systematics on the flux limit for each source. Table 3 shows the neutrino flux limit for each source. Figure 4 shows, by declination, the neutrino flux limit and the neutrino flux predicted from the best gamma-ray fit assuming all the gamma-ray emission is from hadronic interaction.
@@
| Name | RA | Dec | Extension | Neutrino 90% CL flux limit | Predicted neutrino flux | IceCube 90% energy range |
| [∘] | [∘] | [∘] | [TeV-1cm-2s-1] | [TeV-1cm-2s-1] | [TeV] | |
| 3HWC J0534+220 | 83.64 | 22.01 | PS | 0.3 - 21.6 | ||
| 3HWC J0634+067 | 98.59 | 6.66 | PS | 0.4 - 143.0 | ||
| 3HWC J1757240 | 269.32 | 23.86 | 0.10 | 98.7 - 6870.0 | ||
| 3HWC J1809190 | 272.42 | 19.34 | 0.23 | 38.7 - 468.3 | ||
| 3HWC J1813125 | 273.40 | 12.70 | 0.27 | 41.3 - 3897.3 | ||
| 3HWC J1813174 | 273.40 | 17.74 | 0.26 | 9.60 - 263.5 | ||
| 3HWC J1825134 | 276.49 | 13.63 | 0.46 | 26.2 - 519.7 | ||
| 3HWC J1831095 | 277.85 | 9.85 | 1.14 | 3.1 - 229.4 | ||
| 3HWC J1837066 | 279.38 | 6.83 | 0.65 | 0.7 - 94.1 | ||
| 3HWC J1843034 | 280.95 | 3.39 | 0.60 | 0.7 - 57.0 | ||
| 3HWC J1847017 | 282.07 | 1.78 | 0.75 | 0.5 - 49.1 | ||
| 3HWC J1849+001 | 282.31 | 0.02 | 0.68 | 0.7 - 56.9 | ||
| 3HWC J1857+027 | 284.35 | 2.82 | 0.60 | 0.4 - 29.2 | ||
| 3HWC J1908+063 | 287.02 | 6.35 | 0.57 | 0.6 - 65.1 | ||
| 3HWC J1912+103 | 288.10 | 10.32 | 0.52 | 0.4 - 24.1 | ||
| 3HWC J1914+118 | 288.71 | 11.79 | 0.71 | 0.3 - 73.9 | ||
| 3HWC J1922+140 | 290.74 | 14.06 | 0.13 | 0.3 - 51.8 | ||
| 3HWC J1928+178 | 292.13 | 17.81 | 0.83 | 0.5 - 42.7 | ||
| 3HWC J1930+188 | 292.56 | 18.81 | 0.76 | 0.5 - 36.6 | ||
| 3HWC J2019+367 | 304.90 | 36.77 | 0.32 | 1.1 - 34.2 | ||
| 3HWC J2031+415 | 308.01 | 41.49 | 0.51 | 0.5 - 25.8 | ||
| 3HWC J2227+610 | 336.82 | 60.94 | PS | 0.4 - 78.4 |
@@
Five sources have a lower neutrino 90% CL flux limit than the predicted neutrino flux. In these cases, we conclude that the gamma-ray emission in the TeV range cannot completely originate from hadronic interactions. If we further assume that the leptonic component shares a similar spectral shape in the TeV energy range, we can constrain the percentage of the gamma-ray flux originating from hadronic interactions. Among these sources, 3HWC J2227+610 was studied by the HAWC collaboration and could be associated with supernova remnant G106.3+2.7. HAWC found a hint for PeV proton acceleration and suggested that 3HWC J2227+610 could be a Galactic PeVatron (Albert et al., 2020a). The Crab Nebula (3HWC J0534+220) is the brightest source in the TeV sky and is the most strongly constrained. 3HWC J1922+140 is close to the W 51 region which contains the star-forming region W51B, the supernova remnant W51C, and a potential PWN, CXOU J192318.5+140305 (Abdalla et al., 2018). 3HWC J1922+140 could be a composite of different astrophysical origins. Table 4 shows the constraints on the hadronic fraction, which range from 51% to 85%.
@@
5 Summary and conclusion
@@
In this study, we present a new method for performing multi-messenger spectral fits with the 3ML software framework (Vianello et al., 2017) using HAWC and IceCube data. We performed a joint fit using HAWC and IceCube data to search for neutrino emission from potential Galactic PeVatrons. No significant neutrino emission was observed. The most significant source from the joint fit is 3HWC J2019+367 with the lowest pre-trial p-value of 0.07, corresponding to a post-trial p-value of 0.21. A binomial test for an excess distributed over several sources was performed. The binomial test gave a minimum p-value of 0.09, obtained for the six most significant sources. This corresponds to a post-trial p-value of 0.34. After trial correction for the two searches, the overall post-trial p-value is 0.37. We calculated upper limits at 90% confidence level on the neutrino flux, assuming the spectral form corresponding to the best fit for gamma rays. We found that five sources have a neutrino flux limit that is lower than the predicted neutrino flux of a purely hadronic source. We conclude that these five sources cannot be purely hadronic and constrain the hadronic fraction of these sources. Next-generation neutrino detectors can improve our understanding of the physical processes of these TeV gamma-ray emitters. The proposed IceCube-Gen2 detector (Aartsen et al., 2021), with an increased effective area compared to IceCube, is expected to become up to five times more sensitive to galactic neutrino sources. Several neutrino experiments, such as the proposed Pacific Ocean Neutrino Experiment (P-ONE) (Agostini et al., 2020), Tropical Deep-sea Neutrino Telescope (TRIDENT) (Ye et al., 2023), KM3NeT (currently under construction) (Margiotta, 2014) and Baikal-GVD (operating and undergoing upgrade) (Avrorin et al., 2022), are located in the northern hemisphere and have the potential to detect galactic neutrino sources near the galactic center. The proposed SWGO experiment, which is a ground-based gamma-ray water Cherenkov observatory in the southern hemisphere, would be able to discover more TeV gamma rays sources in the southern sky and candidates of Galactic PeVatrons (Huentemeyer et al., 2019).
@@ Appendix A Likelihood formalism
Appendix B Pre-trial p-values of the joint search
Table 5 shows the pre-trial p-value of each source.
| Name | pre-trial p-value | post-trial p-value |
| 3HWC J0534+220 | 0.36 | - |
| 3HWC J0634+067 | 0.19 | - |
| 3HWC J1757-240 | 1 | - |
| 3HWC J1809-190 | 1 | - |
| 3HWC J1813-125 | 1 | - |
| 3HWC J1813-174 | 0.70 | - |
| 3HWC J1825-134 | 1 | - |
| 3HWC J1831-095 | 1 | - |
| 3HWC J1837-066 | 1 | - |
| 3HWC J1843-034 | 1 | - |
| 3HWC J1847-017 | 1 | - |
| 3HWC J1849+001 | 1 | - |
| 3HWC J1857+027 | 0.13 | - |
| 3HWC J1908+063 | 0.10 | - |
| 3HWC J1912+103 | 0.08 | - |
| 3HWC J1914+118 | 1 | - |
| 3HWC J1922+140 | 1 | - |
| 3HWC J1928+178 | 0.14 | - |
| 3HWC J1930+188 | 1 | - |
| 3HWC J2019+367 | 0.07 | 0.21 |
| 3HWC J2031+415 | 0.08 | - |
| 3HWC J2227+610 | 1 | - |
Appendix C IceCube Systematics

### v2

@@ 4.1 HAWC best-fit result
First, we present the best-fit results for the gamma-ray emission from the HAWC Pass 5 data alone. Table 2 summarizes the best-fit location, extension, and spectral parameters of each source in our list. In addition, we report the statistical and systematic uncertainty of the spectral fit results. Systematic uncertainties of the HAWC detector include PMT efficiency, charge uncertainty, etc., and are described in Abeysekara et al. (2019b). To evaluate their effect, we perform the model selection using samples simulated with different settings and compare the results with the baseline simulation. The errors of each type of systematic uncertainty are added in quadrature to obtain the total systematic uncertainty, which is given in the table 2. We also reported the energy range of each source. The upper end of this range is defined such that if the spectrum is cut off above this energy the significance of the source drops by . The corresponding definition is used for the lower end of the sensitivity range. The threshold is chosen to reasonably claim that HAWC data can probe the quoted energy range at a significance of or at the 95% confidence level.
Name | RA | Dec | Ext | Nearest TeVCat | Energy range | |||
[∘] | [∘] | [∘] | [TeV-1cm-2s-1] | [TeV] | ||||
3HWC J0534+220 | 83.64 | 22.01 | PS | 2.82 | 0.12 | Crab | 0.1 - 106.1 | |
3HWC J0634+067 | 98.59 | 6.66 | 0.49 | 2.57 | 0 | HAWC J0635+070 | 1.7-53.1 | |
3HWC J1757240 | 269.32 | 23.86 | PS | 2.58 | 0 | HESS J1800240B | 16.9 - 59.7 | |
3HWC J1809190 | 272.42 | 19.34 | 0.23 | 2.09 | 0.23 | HESS J1809193 | 3.0 - 94.5 | |
3HWC J1813125 | 273.4 | 12.7 | 0.27 | 2.6 | 0 | HESS J1813-126 | 2.4 - 59.6 | |
3HWC J1813174 | 273.4 | 17.74 | 0.26 | 1.91 | 0.35 | 2HWC J1814173 | 3.4 - 106.0 | |
3HWC J1825134 | 276.49 | 13.63 | 0.46 | 2.45 | 0.13 | 2HWC J1825134 | 0.5 - 106.0 | |
3HWC J1831095 | 277.85 | 9.85 | 1.14 | 2.59 | 0.14 | HESS J1831098 | 0.4 - 118.7 | |
3HWC J1837066 | 279.38 | 6.83 | 0.65 | 2.69 | 0.1 | 2HWC J1837065 | 0.2 - 133.2 | |
3HWC J1843034 | 280.95 | 3.39 | 0.6 | 2.51 | 0.13 | 2HWC J1844032 | 0.3 - 75.0 | |
3HWC J1847017 | 282.07 | 1.78 | 0.75 | 2.66 | 0.09 | HESS J1848018 | 0.3 - 84.0 | |
3HWC J1849+001 | 282.31 | 0.02 | 0.68 | 2.42 | 0.13 | IGR J184900000 | 0.4 - 149.8 | |
3HWC J1857+027 | 284.35 | 2.82 | 0.6 | 2.73 | 0.13 | HESS J1857+026 | 0.3 - 133.2 | |
3HWC J1908+063 | 287.02 | 6.35 | 0.57 | 2.44 | 0.09 | MGRO J1908+06 | 0.4 - 168.0 | |
3HWC J1912+103 | 288.1 | 10.32 | 0.52 | 2.83 | 0.14 | HESS J1912+101 | 0.4 - 94.4 | |
3HWC J1914+118 | 288.71 | 11.79 | 0.71 | 2.69 | 0 | 2HWC J1914+117* | 0.3 - 106.0 | |
3HWC J1922+140 | 290.74 | 14.06 | 0.13 | 2.77 | 0 | W 51 | 0.2 - 47.3 | |
3HWC J1928+178 | 292.13 | 17.81 | 0.83 | 2.52 | 0.11 | 2HWC J1928+177 | 0.6 - 106.0 | |
3HWC J1930+188 | 292.56 | 18.81 | 0.76 | 2.51 | 0.14 | SNR G054.1+00.3 | 0.7 - 105.9 | |
3HWC J2019+367 | 304.9 | 36.77 | 0.32 | 2.04 | 0.31 | VER J2019+368 | 0.9 - 106.0 | |
3HWC J2031+415 | 308.01 | 41.49 | 0.51 | 2.52 | 0.19 | TeV J2032+4130 | 0.5 - 105.9 | |
3HWC J2227+610 | 336.82 | 60.94 | PS | 2.42 | 0 | Boomerang | 16.7 - 33.5 |
@@ 4.2 Joint-fit result
We performed the joint fit using the best-fit model of HAWC and added a neutrino source as described in Section 3.4. We found no significant neutrino emission. The smallest pre-trial p-value from the individual source search is 0.07 for 3HWC J2019+367, which corresponds to a post-trial p-value of 0.21 when accounting for the trials in the individual source search. We performed the binomial test and found no significant neutrino emission from any subset of the 22 sources. The smallest pre-trial p-value of the binomial test is 0.09 at =6 (3HWC J2019+367, 3HWC J2031+415, 3HWC J1912+103, 3HWC J1908+063, 3HWC J1857+027, 3HWC J1928+178) and corresponds to a post-trial p-value of 0.34 when only accounting for trials in the binomial test. Accounting for both the trials from the individual source search and the binomial test, the final post-trial p-value is 0.37. We also computed the 90% neutrino energy range for each source using the best-fit morphology and spectrum. The 90% neutrino energy range is the central 90% range of true neutrino energy when injecting signal neutrinos with the source’s best-fit gamma-ray morphology and spectrum. IceCube applies a higher energy threshold for neutrino events from the southern sky due to the containment of atmospheric muon, resulting in a higher energy range for southern sources.
@@
Based on the null result of the joint search, we also calculated the neutrino flux limits at 90% CL for the 22 sources (see table 3). We account for IceCube systematics by calculating the sensitivity using different Monte Carlo systematic sets and adding the difference to the baseline in quadrature. Appendix B details the effect of IceCube systematics on the flux limit for each source. Table 3 shows the neutrino flux limit for each source. Figure 4 shows, by declination, the neutrino flux limit and the neutrino flux predicted from the best gamma-ray fit assuming all the gamma-ray emission is from hadronic interaction.
@@
Name | RA | Dec | Extension | p-value | Neutrino 90% CL flux limit | Predicted neutrino flux | Energy range |
[∘] | [∘] | [∘] | [TeV-1cm-2s-1] | [TeV-1cm-2s-1] | [TeV] | ||
3HWC J0534+220 | 83.64 | 22.01 | PS | 0.36 | 0.3 - 21.6 | ||
3HWC J0634+067 | 98.59 | 6.66 | PS | 0.19 | 0.4 - 143.0 | ||
3HWC J1757240 | 269.32 | 23.86 | 0.10 | 1 | 98.7 - 6870.0 | ||
3HWC J1809190 | 272.42 | 19.34 | 0.23 | 1 | 38.7 - 468.3 | ||
3HWC J1813125 | 273.40 | 12.70 | 0.27 | 1 | 41.3 - 3897.3 | ||
3HWC J1813174 | 273.40 | 17.74 | 0.26 | 0.7 | 9.60 - 263.5 | ||
3HWC J1825134 | 276.49 | 13.63 | 0.46 | 1 | 26.2 - 519.7 | ||
3HWC J1831095 | 277.85 | 9.85 | 1.14 | 1 | 3.1 - 229.4 | ||
3HWC J1837066 | 279.38 | 6.83 | 0.65 | 1 | 0.7 - 94.1 | ||
3HWC J1843034 | 280.95 | 3.39 | 0.60 | 1 | 0.7 - 57.0 | ||
3HWC J1847017 | 282.07 | 1.78 | 0.75 | 1 | 0.5 - 49.1 | ||
3HWC J1849+001 | 282.31 | 0.02 | 0.68 | 1 | 0.7 - 56.9 | ||
3HWC J1857+027 | 284.35 | 2.82 | 0.60 | 0.13 | 0.4 - 29.2 | ||
3HWC J1908+063 | 287.02 | 6.35 | 0.57 | 0.1 | 0.6 - 65.1 | ||
3HWC J1912+103 | 288.10 | 10.32 | 0.52 | 0.08 | 0.4 - 24.1 | ||
3HWC J1914+118 | 288.71 | 11.79 | 0.71 | 1 | 0.3 - 73.9 | ||
3HWC J1922+140 | 290.74 | 14.06 | 0.13 | 1 | 0.3 - 51.8 | ||
3HWC J1928+178 | 292.13 | 17.81 | 0.83 | 0.14 | 0.5 - 42.7 | ||
3HWC J1930+188 | 292.56 | 18.81 | 0.76 | 0.1 | 0.5 - 36.6 | ||
3HWC J2019+367 | 304.90 | 36.77 | 0.32 | 0.07 | 1.1 - 34.2 | ||
3HWC J2031+415 | 308.01 | 41.49 | 0.51 | 0.08 | 0.5 - 25.8 | ||
3HWC J2227+610 | 336.82 | 60.94 | PS | 1 | 0.4 - 78.4 |
@@
Five sources have a lower neutrino 90% CL flux limit than the predicted neutrino flux. In these cases, we conclude that the gamma-ray emission detected in the HAWC energy range cannot completely originate from hadronic interactions. This assumes that the HAWC observed spectrum extends to the IceCube energy range and it doesn’t account for HAWC spectral uncertainty. If we further assume that the leptonic component shares a similar spectral shape in the TeV energy range, we can constrain the percentage of the gamma-ray flux originating from hadronic interactions. Among these sources, 3HWC J2227+610 was studied by the HAWC collaboration and could be associated with supernova remnant G106.3+2.7. HAWC found a hint for PeV proton acceleration and suggested that 3HWC J2227+610 could be a Galactic PeVatron (Albert et al., 2020a). The Crab Nebula (3HWC J0534+220) is the brightest source in the TeV sky and is the most strongly constrained. However, the Crab Nebula is widely believed to be a leptonic source. A more sensitive neutrino instrument is required to put a relevant constraint, in the 1% range, on the hadronic fraction. 3HWC J1922+140 is close to the W 51 region which contains the star-forming region W51B, the supernova remnant W51C, and a potential PWN, CXOU J192318.5+140305 (Abdalla et al., 2018). 3HWC J1922+140 could be a composite of different astrophysical origins. Table 4 shows the constraints on the hadronic fraction, which range from 51% to 85%.
@@
5 Summary and discussion
In this paper, we present a new method for performing multi-messenger spectral fits with the 3ML software framework (Vianello et al., 2017) using HAWC and IceCube data. This is motivated by the predicted neutrino emission by Galactic cosmic-ray accelerators with gamma-ray emission. The new joint-fitting method with gamma-ray data and neutrino data brings improvement in search sensitivity due to more detailed modeling of the neutrino emission under the assumption that the gamma rays and neutrinos originated from the same population of hadrons. In addition, we could provide a direct constraint on the hadronic ratio in the case of null detection using the emission model derived from the gamma-ray data if the assumption is true. We stress the caveat that a mis-modeling of the emission region, for example, due to gamma rays and neutrinos being produced in different regions of a source, could weaken the constraints presented in this paper.
@@ 5 Summary and discussion
We used the HAWC data to obtain the best-fit model for the gamma-ray emission. We performed a joint fit using HAWC and IceCube data to search for neutrino emission from potential Galactic cosmic-ray accelerators. No significant neutrino emission was observed. The most significant source from the joint fit is 3HWC J2019+367 with the lowest pre-trial p-value of 0.07, corresponding to a post-trial p-value of 0.21. A binomial test for an excess distributed over several sources was performed. The binomial test gave a pre-trial p-value of 0.09, obtained for the six most significant sources and corresponds to a post-trial p-value of 0.34. After trial correction for the two searches, the overall post-trial p-value is 0.37. We calculated upper limits at 90% confidence level on the neutrino flux, assuming the spectral form corresponding to the best fit for gamma rays. We found that five sources have a neutrino flux limit that is lower than the predicted neutrino flux of a purely hadronic source. We conclude that these five sources cannot be purely hadronic and constrain the hadronic fraction of these sources.
We found that for some of the sources, the HAWC data do not show a 2 significance for emission above 100 TeV, meaning that the analysis would not warrant the conclusion that those sources are PeVatrons even if neutrinos are detected from the source. In addition, the neutrino 90% energy range for some of the sources does not exceed 50 TeV, meaning the neutrino data would likely not be grounds to establish PeV cosmic-ray acceleration in the case of detection. A more sensitive experiment such as LHAASO could potentially determine the limit on the maximum energy of the source, and neutrino detection could act as evidence of hadronic interactions. In addition, it could provide more PeVatron candidates with better gamma-ray spectrum and morphology measurement. The recently released 1LHAASO catalog contains 43 sources with gamma-ray emission above 100 TeV (Cao et al., 2024b) and a joint search with neutrino data will provide better sensitivity and constraints on their hadronic emission. Next-generation neutrino detectors can improve our understanding of the physical processes of these TeV gamma-ray emitters. The proposed IceCube-Gen2 detector (Aartsen et al., 2021), with an increased effective area compared to IceCube, is expected to become up to five times more sensitive to galactic neutrino sources. Several neutrino experiments, such as the proposed Pacific Ocean Neutrino Experiment (P-ONE) (Agostini et al., 2020), Tropical Deep-sea Neutrino Telescope (TRIDENT) (Ye et al., 2023), KM3NeT (currently under construction) (Margiotta, 2014) and Baikal-GVD (operating and undergoing upgrade) (Avrorin et al., 2022), are located in the northern hemisphere and have the potential to detect galactic neutrino sources near the galactic center. The proposed SWGO experiment, which is a ground-based gamma-ray water Cherenkov observatory in the southern hemisphere, would be able to discover more TeV gamma rays sources in the southern sky and candidates of Galactic PeVatrons (Huentemeyer et al., 2019).
@@ Appendix A Likelihood formalism
Appendix B IceCube Systematics